Nicole:

Determine if the lines that pass through the given points are parallel, perpendicular or neither. Line A: (5, 10) and (12, 6) Line B: (-2, 4) and (5, 8)

Nicole:

@dude

dude:

For this you just need to find the slope between the points \(m_a=\dfrac{y_2-y_1}{x_2-x_1}\) and \(m_b=\dfrac{y_2-y_1}{x_2-x_1}\)

Nicole:

ma:-4/7 mb: 4/7

Nicole:

now what do I do

Nicole:

@dude

Nicole:

@Gdeinward

Gdeinward:

the slopes indicate that the points are located on opposite sides of the y-axis

Gdeinward:

Idealy, if you cant visualize it, you would graph the points to determine their relationship.

Nicole:

so parallel?

Nicole:

@Gdeinward am i correct?

Gdeinward:

no. Here is a grpah with the points:

Gdeinward:

1 attachment
Nicole:

but I learned that we are supposed to find the slope and base it off that

Gdeinward:

There are general rules that decide wether they ar eparralel or night. They are as follows:

Gdeinward:

perpendicular slopes are negative reciprocals of each other

Gdeinward:

the slopes of parallel lines are equal.

Gdeinward:

So if the slopes dont reciprocate any of these rules, then it must be neither.

Nicole:

so -4/7 and 4/7 are not parallel or perpendicular?

dude:

Yep, theyre neither They just intersect

Gdeinward:

That is correct.

Gdeinward:

Like Dude said.

Nicole:

oh okay and this one: Determine if the lines that pass through the given points are parallel, perpendicular or neither. Line A: (2, 3) and (-4, 5) Line B: (-4, 9) and (-3, 12)

Nicole:

the slopes are 2/-6 and 3/1

Gdeinward:

Well, following the previous rules, what do you think?

Nicole:

perpendicular

Nicole:

when its simplified

Gdeinward:

That is correct. Now ur getting the hang of it.

Nicole:

1 attachment
Nicole:

@dude

dude:

Make sure to make a new post Anyway, well this looks like a standard equation Do you know the slope of that line is?